### v6

@@ Abstract
1 Introduction
In this paper, we are concerned with the following minimization problem with $m$ equality constraints,
@@ 1 Introduction
Traditional algorithms like the Lagrange Multipliers [38] and the Alternating Direction Method of Multipliers (ADMM) [43] lack guarantees of converging to the global constrained minimizer when dealing with non-convex or non-differentiable loss functions $\mathcal{E}(v)$. A new framework is required to effectively handle such cases, and recently, a class of gradient-free methods called consensus-based optimization (CBO) methods [4, 11, 35, 41] have emerged as promising approaches for handling non-convex and non-differentiable loss functions. Motivated by the well-known Laplace’s principle [6, 16, 33], they are decentralized and gradient-free algorithms that leverage the power of information sharing and cooperation among individual particles. However, it is important to highlight that much of the existing work has focused on the unconstrained case such as [11, 12, 15, 22, 23, 24, 28, 29, 30, 31, 36, 37, 42]. We refer the readers to survey articles [13, 27, 40] for a more detailed and complete summary. For the efficient implementation of CBO, we refer to the CBX packages developed for Python and Julia [4].
@@
Limited work has been done for the constrained case. The primary challenge lies in reconciling the CBO model’s tendency to drive agents towards the global minimizer with the need for agents to remain within the constraint set and converge to the constrained minimizer. Currently, there are mainly two approaches. One involves projection onto the hypersurface [2, 18, 19, 20]. However, this method requires computing the distance function $\text{dist}(\Gamma,v)=\inf\{\|v-u\|_{2}\mid u\in\Gamma\}$ with $\Gamma$ representing the constraint set. Extending this method to handle general multiple equality constraints is not straightforward. In cases where the constraint set $\Gamma$ is complicated, this computation of $\text{dist}(\Gamma,v)$ becomes infeasible. Another method introduces constraints as a penalization term in the objective function [7, 13], transforming it into an unconstrained problem for CBO. However, the convergence is sensitive to the landscape of the objective function and the penalization constant, which makes it difficult to achieve high accuracy.
@@
In this paper, we introduce a third strategy for constrained CBO along with convergence analysis and numerical experiments. Instead of performing projection onto the constraint set or adding penalty terms, we propose a novel approach that combines the classical unconstrained CBO algorithm with gradient descent on the function $G(v)=\sum_{i=1}^{m}g^{2}_{i}(v)$, serving as a forcing term to the constraint set. Importantly, we do not require the differentiability of the target function $\mathcal{E}$ and only need a mild differentiability condition on $G$. Compared with the other two constrained CBO methods, our method applies to general equality constraints, achieves faster convergence, and has consistently more stable performance as shown in Figures 1 and 2.
1.1 Contributions
@@
Our main contributions are threefold. Firstly, we introduce a new CBO-based method for solving constrained optimization problems, with possibly non-convex and non-differentiable objective functions. This method can accommodate a wide range of equality constraints, including the ability to handle multiple constraints concurrently. Secondly, we provide rigorous theoretical guarantees for the continuous-in-time model of the proposed method. Specifically, we establish the mean-field limit of the method and provide a detailed analysis of its convergence behavior within this limit, using a new quantitative Laplace Principle that differs from the unconstrained version in [22]. This principle enables us to quantify the contributions of the consensus dynamics and gradient descent, providing a framework for handling constraint sets. See Remark 4.8 for the details. Thirdly, we present a stable discretized algorithm designed to approximate the dynamics of the continuous-in-time model efficiently. Notably, this algorithm handles the stiff term of order $O(\epsilon^{-1})$ without requiring the time step to approach zero when $\epsilon$ becomes very small.
1.2 Organizations
@@ 1.3 Notations
2 The dynamics of the constrained consensus-based optimization algorithm
@@
(1) $\min_{v\in\mathbb{R}^{d}}\ \mathcal{E}(v)\quad\text{s.t.}\quad g_{i}(v)=0,\ i=1,\ldots,m.$
Here, we require the function $g_{i}(x)$ to be first-order differentiable. It is noteworthy that Problem (1) can be reformulated equivalently as follows:
@@
where $G(v)=\sum_{i=1}^{m}g^{2}_{i}(v)$. Our method will be based on formulation (2).
@@
(3) $dV^{i,N}_{t}=-\lambda\left(V^{i,N}_{t}-v_{\alpha}\left(\hat{\rho}_{t}^{N}\right)\right)\,dt-\dfrac{1}{\epsilon}\nabla G\left(V^{i,N}_{t}\right)\,dt+\sigma D^{i,N}_{t}\,dB^{i,N}_{t},\quad V^{i,N}_{0}\sim\rho_{0},$
where
@@
The diffusion term $\sigma D^{i,N}_{t}\,dB^{i,N}_{t}$ encourages particles to explore the landscape of $\mathcal{E}(v)$, where $D^{i,N}_{t}$ is a $d\times d$ matrix function that determines the way in which particles explore the landscape and $\{B_{t}^{i,N}\}_{i=1,...,N}$ are independent Wiener processes. There are different choices for the matrix function $D^{i,N}_{t}$ [35, 12]. In this paper, we use the anisotropic exploration defined as,
@@
The third term $\tfrac{1}{\epsilon}\nabla G\big{(}V^{i,N}_{t}\big{)}\,dt,$ addresses the constraint $\{G=0\}$. Since 0 is the minimum of the non-negative function $G(v)$, finding the constraint $\{G(v)=0\}$ is the same as minimizing $G(v)$. Therefore, we propose the third term as a gradient descent of $G(v)$, allowing $G(v)$ to be minimized during the algorithm’s progression. Here $\epsilon>0$ is a parameter that controls the magnitude of this term. When $\epsilon$ is small, this term will encourage particles to concentrate around $\{G=0\}$. These ideas were used in kinetic equations for swarming including alignment terms of Cucker-Smale type in order to derive kinetic models on the sphere such as the Viczek-Fokker-Planck model, see [8, 9, 10, 1] for instance.
Before we proceed to the theoretical analysis of the model, we first present a comparison result in Figures 1 and 2 to illustrate the superior performance of the proposed interacting particle system (3) compared to the projected CBO system [18] and the penalized CBO system [7] on a two-dimensional Ackley function (shown in Figure 1) with different constraints.
We defer algorithmic formulation to Section 5, and details of the experiments to Appendix K.1, respectively. Our method achieves a $100\%$ success rate in finding the unconstrained minimizer and demonstrates the fastest convergence rate in all experiments. The projected CBO performs similarly to our method when the constraint is a circle, but it is not applicable to parabolic curves. In contrast, the Penalized CBO exhibits a significantly lower success rate due to two main reasons: First, when the constrained minimizer is not a local minimizer of the objective function, the global minimizer $v_{p}^{*}$ of the penalized objective function usually differs from the constrained minimizer $v^{*}$. Second, although it is possible to increase the penalty sufficiently to reduce the distance $\|v_{p}^{*}-v^{*}\|_{2}$, the landscape is dominated by the penalized term, making the objective function resemble a minor perturbation around the penalty. Consequently, it becomes more challenging for the optimization method to locate the global minimizer, and leads to a longer time for CBO to converge.
@@
Figure 1: Objective function and success rate of three constrained CBO methods.
@@
Figure 2: The averaged distance to the true constrained minimizer over $100$ simulations.
3 Well-posedness and Mean-field limit
@@
Throughout this section, we make the following assumptions. {assumption} (1) The function $\mathcal{E}$ is bounded with $\inf\mathcal{E}=\underline{\mathcal{E}}$ and $\sup\mathcal{E}=\overline{\mathcal{E}}$.
@@
Briefly speaking, in Assumption 3 (1) and (2), we assume the loss function $\mathcal{E}$ is bounded, locally Lipschitz and with at most quadratic growth. In Assumption 3 (3) and (4), we assume the gradient of the function $G$ is globally Lipschitz and with at most linear growth. We establish the well-posedness of the interacting particle system (6) in the following theorem.
@@
$\frac{d}{dt}\langle\phi,\rho_{t}\rangle$ $=-\lambda\langle(v-v_{\alpha}\left(\rho_{t}\right))\cdot\nabla\phi,\rho_{t}\rangle-\frac{1}{\epsilon}\langle\nabla G\cdot\nabla\phi,\rho_{t}\rangle$
$\quad+\frac{\sigma^{2}}{2}\sum_{k=1}^{d}\langle(v-v_{\alpha}\left(\rho_{t}\right))_{k}^{2}\partial_{kk}\phi,\rho_{t}\rangle,$
Remark 3.3.
In the Definition 3.2 (ii), the test function space is $\mathcal{C}^{2}_{c}(\mathbb{R}^{d})$. We could extend $\mathcal{C}^{2}_{c}(\mathbb{R}^{d})$ to a larger space $\mathcal{C}^{2}_{*}(\mathbb{R}^{d})$ as explained in Appendix C, which will be used in the proof later. $\mathcal{C}^{2}_{*}(\mathbb{R}^{d})$ is defined below.
$\mathcal{C}^{2}_{*}(\mathbb{R}^{d}):=\big{\{}\phi\in\mathcal{C}^{2}(\mathbb{R}^{d})|\text{ }|\partial_{k}\phi(x)|\leq C(1+|x_{k}|)\text{ and }\sup_{x\in\mathbb{R}^{d}}|\partial_{kk}\phi(x)|<\infty\text{ for all $k=1,2,...,d$}\big{\}}.$
@@
Then we present the result showing that (7), (8) indeed characterize the mean-field limit of the particle system.
@@
4 Convergence to the constrained minimizer in the mean-field limit
In this section, we will analyze the behavior of the weak solution of the Fokker-Planck equation (8). Throughout this section, we assume Problem (2) admits a unique solution $v^{*}$, which is a common assumption in both unconstrained and constrained CBO-related literature, e.g., [7, 12, 19, 22] [^1] In [7], the uniqueness assumption is slightly weaker: it requires the penalized loss function to admit a unique global minimizer, which coincides with the constrained minimizer.. Our primary goal is to establish a key result: under suitable assumptions and the selection of appropriate parameters, the particles will concentrate around $v^{*}$ with arbitrary closeness, confirming the effectiveness of the method in the mean-field limit.
@@ 4.1 Main Results
Suppose $G$ and $\mathcal{E}$ satisfy Assumption 4.2 (well-behaved). Fix any $\tau\in(0,1)$ and parameters $\lambda,\sigma>0$ with $2\lambda>\sigma^{2}$. There exists a function $I:\mathbb{R}\rightarrow\mathbb{R}$ such that for any error tolerance $\delta\in\big{(}0,\mathcal{V}(\rho_{0})\big{)}$, as long as $\rho_{0}\big{(}B(v^{*},r)\big{)}>0$ for all $r>0$ and $\int G\,d\rho_{0}(v)\leq I(\delta)$, then one can find $\alpha$ and $\epsilon$ so that
(10) $\min_{t\in[0,T^{*}]}\mathcal{V}\left(\rho_{t}\right)\leq\delta,\quad\text{where}\quad T^{*}=\dfrac{1}{(1-\tau)(2\lambda-\sigma^{2})}\log\left(\tfrac{\mathcal{V}\left(\rho_{0}\right)}{\delta}\right).$
Furthermore, until $\mathcal{V}\left(\rho_{t}\right)$ reaches the prescribed accuracy $\delta$, the following exponential decay holds:
(11) $\mathcal{V}\left(\rho_{t}\right)\leq\mathcal{V}(\rho_{0})\exp\big{(}-(1-\tau)(2\lambda-\sigma^{2})t\big{)}.$
Remark 4.2.
@@
4.2 Assumption
In this subsection, we define clearly what it means by being ‘well-behaved’. $G$ and $\mathcal{E}$ are well-behaved if the following Assumption 4.2 is satisfied. It is worth noting that Assumption 4.2 in this section is independent of Assumption 3. In other words, for the proofs in this section, Assumption 3 is not required.
{assumption}

### v8

@@ Abstract.
1 Mathematical Institute, University of Oxford, United Kingdom. (carrillo@maths.ox.ac.uk)
2 School of Mathematical Sciences, Institute of Natural Sciences, MOE-LSC, Shanghai Jiao Tong University, Shanghai, 200240, P. R. China. (shijin-m@sjtu.edu.cn)
3 Department of Mathematics, University of California-San Diego, La Jolla, California 92093, USA. (haz053@ucsd.edu)
4 Department of Statistics and Data Science, University of California, Los Angeles, USA (yuz244@ucsd.edu)
All authors contributed equally to this work and are listed alphabetically.
$\dagger$ Corresponding author.
1. Introduction
In this paper, we are concerned with the following minimization problem with equality constraint,
@@ 1. Introduction
Traditional algorithms like the Lagrange Multipliers [46] and the Alternating Direction Method of Multipliers (ADMM) [51] lack guarantees of converging to the global constrained minimizer when dealing with non-convex or non-differentiable loss functions $\mathcal{E}(v)$. A new framework is required to effectively handle such cases, and recently, a class of gradient-free methods called consensus-based optimization (CBO) methods [4, 12, 41, 48] have emerged as promising approaches for handling non-convex and non-differentiable loss functions. Motivated by the well-known Laplace’s principle [7, 17, 38], they are decentralized and gradient-free algorithms that leverage the power of information sharing and cooperation among individual particles. However, it is important to highlight that much of the existing work has focused on the unconstrained case, see [12, 13, 16, 23, 24, 25, 29, 30, 31, 32, 44, 43, 49]. We refer the readers to survey articles [14, 28, 50] for a more detailed and complete summary. For the efficient implementation of CBO, we refer to the CBX packages developed for Python and Julia [4].
@@
Limited work has been done for the constrained case. The primary challenge lies in reconciling the CBO model’s tendency to drive agents towards the global minimizer with the need for agents to remain within the constraint set and converge to the constrained minimizer. Currently, there are mainly two approaches. One involves projection onto the hypersurface [2, 19, 20, 21]. However, this method requires computing the derivatives of the distance function $\operatorname{dist}(\Gamma,v)=\inf\{\|v-u\|_{2}\mid u\in\Gamma\}$ with $\Gamma$ representing the constraint set. Extending this method to handle general multiple equality constraints is not straightforward. In cases where the constraint set $\Gamma$ is complicated, this computation about $\operatorname{dist}(\Gamma,v)$ becomes infeasible. Another method introduces constraints as a penalization term in the objective function [8, 14], transforming it into an unconstrained problem for CBO. However, the convergence is sensitive to the landscape of the objective function and the penalization constant, which makes it difficult to achieve high accuracy.
@@
In this paper, we introduce a third strategy for constrained CBO along with convergence analysis and numerical experiments. Instead of performing projection onto the constraint set or adding penalty terms, we propose a novel approach that combines the classical unconstrained CBO algorithm with gradient descent on the function $G(v)={g^{2}(v)}$, serving as a forcing term to the constraint set. Importantly, we do not require the differentiability of the target function $\mathcal{E}$ and only need a mild differentiability condition on $G$. Compared with the other two constrained CBO methods (projected CBO and penalized CBO), our method applies to general equality constraints, achieves faster convergence, and has consistently more stable performance as shown in Figures 1 and 2.
1.1. Contributions
@@
Our main contributions are threefold. Firstly, we introduce a new CBO-based method for solving constrained optimization problems, with possibly non-convex and non-differentiable objective functions. This method can accommodate a wide range of equality constraints, including the ability to handle multiple constraints concurrently. Secondly, we provide rigorous theoretical guarantees for the continuous-in-time model of the proposed method. Specifically, we establish the mean-field limit of the method and provide a detailed analysis of its convergence behavior within this limit, using a new quantitative Laplace Principle that differs from the unconstrained version in [23]. This principle enables us to quantify the contributions of the consensus dynamics and gradient descent, providing a framework for handling constraint sets. See Remark 5 for the details. Thirdly, we present a stable discretized algorithm designed to approximate the dynamics of the continuous-in-time model efficiently. Notably, this algorithm handles the stiff term of order $O(\epsilon^{-1})$ without requiring the time step to approach zero when $\epsilon$ becomes very small.
1.2. Organizations
@@ 1.3. Notations
2. The dynamics of the constrained consensus-based optimization algorithm
@@
Here, we require the function $g(x)$ to be first-order differentiable. It is noteworthy that Problem (1) can be reformulated equivalently as follows:
where $G(v)={g^{2}(v)}$. Our method will be based on formulation (2).
@@
(3) $\begin{gathered}dV^{i,N}_{t}=-\lambda\left(V^{i,N}_{t}-v_{\alpha}\left(\hat{\rho}_{t}^{N}\right)\right)\,dt-\dfrac{1}{\epsilon}\nabla G\left(V^{i,N}_{t}\right)\,dt+\sigma D^{i,N}_{t}\,dB^{i,N}_{t},\end{gathered}$
with $V^{i,N}_{0}\sim\rho_{0}$, where
@@
The diffusion term $\sigma D^{i,N}_{t}\,dB^{i,N}_{t}$ encourages particles to explore the landscape of $\mathcal{E}(v)$, where $D^{i,N}_{t}$ is a $d\times d$ matrix function that determines the way in which particles explore the landscape and $\{B_{t}^{i,N}\}_{i=1,...,N}$ are independent Wiener processes. There are different choices for the matrix function $D^{i,N}_{t},$ see [41, 13]. In this paper, we use the anisotropic exploration defined as,
@@
The third term $\tfrac{1}{\epsilon}\nabla G\big(V^{i,N}_{t}\big)\,dt,$ addresses the constraint $\{G=0\}$. Since 0 is the minimum of the non-negative function $G(v)$, finding the constraint $\{G(v)=0\}$ is the same as minimizing $G(v)$. Therefore, we propose the third term as a gradient descent of $G(v)$, allowing $G(v)$ to be minimized during the algorithm’s progression. Here $\epsilon>0$ is a parameter that controls the magnitude of this term. When $\epsilon$ is small, this term will encourage particles to concentrate around $\{G=0\}$. These ideas were used in kinetic equations for swarming including alignment terms of Cucker-Smale type in order to derive kinetic models on the sphere such as the Viczek-Fokker-Planck model, see [9, 10, 11, 1] for instance. Inevitably, taking $\epsilon$ small can introduce numerical stiffness and instability, so this term must be handled carefully in the discretization. In Section 5, a stable discrete scheme is derived to treat this term.
Before we proceed to the theoretical analysis of the model, we first present a preliminary numerical result in Figures 1 and 2 to illustrate the superior performance of the proposed interacting particle system (3) compared to the projected CBO system [19] and the penalized CBO system [8] on a two-dimensional Ackley function (shown in Figure 1) with different constraints.
We defer algorithmic formulation to Section 5, and details of the preliminary experiments to Appendix K.1. Our method achieves a $100\%$ success rate in finding the unconstrained minimizer and demonstrates the fastest convergence rate in all three cases as shown in Figure 1 (b). The projected CBO performs similarly to our method when the constraint is a circle, but it is not directly applicable to parabolic curves [^1] For parabolic curve $v_{1}^{2}=v_{2}$, each evaluation involves computing the closest-point projection by solving a cubic optimality condition and selecting the minimizing real root. This per-iteration overhead is considerably higher than for the other approaches, so we omit this experiment in the comparison.. In contrast, the Penalized CBO exhibits a significantly lower success rate due to two main reasons: First, when the constrained minimizer is not a local minimizer of the objective function, the global minimizer $v_{p}^{*}$ of the penalized objective function usually differs from the constrained minimizer $v^{*}$. Second, although it is possible to increase the penalty sufficiently to reduce the distance $\|v_{p}^{*}-v^{*}\|_{2}$, the landscape is dominated by the penalized term, making the objective function resemble a minor perturbation around the penalty. Consequently, it becomes more challenging for the optimization method to locate the global minimizer, and leads to a longer time for CBO to converge.
@@
Figure 1. Objective function and success rate of three constrained CBO methods.
@@
Figure 2. The averaged distance to the true constrained minimizer over $100$ simulations.
3. Well-posedness and Mean-field limit
@@
Throughout this section, we make the following assumptions.
Assumption 1.
(1) The function $\mathcal{E}$ is bounded with $\inf\mathcal{E}=\underline{\mathcal{E}}$ and $\sup\mathcal{E}=\overline{\mathcal{E}}$.
@@
Briefly speaking, in Assumption 1 (1) and (2), we assume the loss function $\mathcal{E}$ is bounded, locally Lipschitz and with at most quadratic growth. In Assumption 1 (3) and (4), we assume the gradient of the function $G$ is globally Lipschitz and with at most linear growth. Now we establish the well-posedness of the interacting particle system (6) in the following theorem.
@@
$\frac{d}{dt}\langle\phi,\rho_{t}\rangle$ $=-\lambda\langle(v-v_{\alpha}\big(\rho_{t}\big))\cdot\nabla\phi,\rho_{t}\rangle-\frac{1}{\epsilon}\langle\nabla G\cdot\nabla\phi,\rho_{t}\rangle$
$\quad+\frac{\sigma^{2}}{2}\sum_{k=1}^{d}\langle(v-v_{\alpha}\big(\rho_{t}\big))_{k}^{2}\partial_{kk}\phi,\rho_{t}\rangle,$
Remark 1.
In the Definition 3.2 (ii), the test function space is $\mathcal{C}^{2}_{c}(\mathbb{R}^{d})$. We could extend $\mathcal{C}^{2}_{c}(\mathbb{R}^{d})$ to a larger space $\mathcal{C}^{2}_{*}(\mathbb{R}^{d})$ as explained in Appendix D, which will be used in the proof later. $\mathcal{C}^{2}_{*}(\mathbb{R}^{d})$ is defined below.
$\mathcal{C}^{2}_{*}(\mathbb{R}^{d}):=\big\{$ $\phi\in\mathcal{C}^{2}(\mathbb{R}^{d})\mid|\partial_{k}\phi(x)|\leq C(1+|x_{k}|)$
$\text{ and }\sup_{x\in\mathbb{R}^{d}}|\partial_{kk}\phi(x)|<\infty\text{ for all $k=1,2,...,d$}\big\}.$
@@
Then we are ready to present the result showing that (7), (8) indeed characterize the mean-field limit of the particle system.
@@
4. Convergence to the constrained minimizer in the mean-field limit
In this section, we will analyze the behavior of the weak solution of the Fokker-Planck equation (8), which charaterizes the mean-field behavior of the proposed finite-particle system. Throughout this section, we assume Problem (2) admits a unique solution $v^{*}$, which is a common assumption in both unconstrained and constrained CBO-related literature, e.g., see [8, 13, 20, 23] [^2] In [8], the uniqueness assumption is slightly weaker: it requires the penalized loss function to admit a unique global minimizer, which coincides with the constrained minimizer.. Our primary goal is to establish a key result: under suitable assumptions and the selection of appropriate parameters, the particles will concentrate around $v^{*}$ with arbitrary closeness, confirming the effectiveness of the method in the mean-field limit.
@@ 4.1. Main Results
Suppose $G$ and $\mathcal{E}$ satisfy Assumption 2 (well-behaved). Fix any $\tau\in(0,1)$ and parameters $\lambda,\sigma>0$ with $2\lambda>\sigma^{2}$. There exists a function $I:\mathbb{R}\rightarrow\mathbb{R}$ such that for any error tolerance $\delta\in\big(0,\mathcal{V}(\rho_{0})\big)$, as long as $\rho_{0}\big(B(v^{*},r)\big)>0$ for all $r>0$ and $\int G\,d\rho_{0}(v)\leq I(\delta)$, then one can find $\alpha$ and $\epsilon$ so that
(10) $\min_{t\in[0,T^{*}]}\mathcal{V}\left(\rho_{t}\right)\leq\delta,\quad\text{where}\quad T^{*}=\dfrac{1}{(1-\tau)(2\lambda-\sigma^{2})}\log\left({\mathcal{V}\left(\rho_{0}\right)}/{\delta}\right).$
Furthermore, until $\mathcal{V}\big(\rho_{t}\big)$ reaches the prescribed accuracy $\delta$, the following exponential decay holds:
(11) $\mathcal{V}\big(\rho_{t}\big)\leq\mathcal{V}(\rho_{0})\exp\big(-(1-\tau)(2\lambda-\sigma^{2})t\big).$
Remark 2.
@@
4.2. Assumption
In this subsection, we define clearly what it means by being ‘well-behaved’. $G$ and $\mathcal{E}$ are well-behaved if the following Assumption 2 is satisfied. It is worth noting that Assumption 2 in this section is independent of Assumption 1. In other words, for the proofs in this section, Assumption 1 is not required.
Assumption 2.